An aeroplane is moving in a circular path at 250 km/hr. What is the change in velocity after half a revolution?


 

Grade:12

2 Answers

vikas askiitian expert
13 years ago

after half revolution direction of velocity is opposite to initial direction ...

if initial direction is along +x axis then after half revolution its direction will become towards -x ..

initial velocity = 250 (i)

final velocity  = 250(-i)

change = 250 + 250 = 500

approve if u like my ans

Shaima Haque
6 years ago
change in velocity=2vsin(o/2)  {let o be the angle of rotation at the centre of the circle}
o is 180(for half a revolution)
so,
change in velocity=2vsin90
=2v
=2*250
=500km/hr